OpenStudy (anonymous):

f(x) = 4x - 7, g(x) = 6x + 8 ... find (fg)(x) (my answer is 24x^2-10x-56 ... does this look right?)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Actually, not quite, f(g(x))=4(6x+8)-7=24x+32-7=24x+25

OpenStudy (anonymous):

or were you looking for f(x)*g(x)?

OpenStudy (anonymous):

the problem is written as find (fg)(x)

OpenStudy (anonymous):

That usually means f(g(x)), but notation does sometimes vary..

OpenStudy (anonymous):

no it doesnt

OpenStudy (anonymous):

f(g(x)) is written like that, f(g(x))

OpenStudy (anonymous):

or I have seen (f o g) (x)

OpenStudy (anonymous):

Actually, you will find that I am right, I am a statistics and economics major at UC Berkeley. Check this article to verify: http://www.ehow.com/how_5203413_fgx.html :)

OpenStudy (anonymous):

wrong noob cake (fg)(x) is the product of two functions notation , you idiot

OpenStudy (anonymous):

and fg (X) would never be written as multiplication would it?

OpenStudy (anonymous):

:| the stupidity and you are meant to be smart

OpenStudy (anonymous):

I know, why don't you troll a bit and throw insults at people randomly because they can't see you or reply to your face. People show no respect on the internet.. we are both trying to help someone, can't you see that and put petty insults aside? I think the notation is unclear: here is another topic where people are discussing the same issue: http://www.physicsforums.com/showthread.php?t=237718

OpenStudy (anonymous):

@techcrew: I would ask your teacher, and maybe put both answers :)

OpenStudy (anonymous):

there is a difference between fg(x) and (fg)(x) , god, are u blind ? mr stats major, do you know what brackets mean?

OpenStudy (anonymous):

jesus, math students are so stupid

OpenStudy (anonymous):

the answers all are polynomials...it is multiple choice. they are all squared so i think it would be fx*gx

OpenStudy (anonymous):

well that simplifies it.

OpenStudy (anonymous):

lols man , confirmed my theory, math students are completely useless :P . Seriously some of guys lack common sense, you can do all your pure maths which has no applications to society but you cant even apply common sense lol

OpenStudy (anonymous):

whoa, you got a math degree, big deal , I am second year engineering student ( probably done like 1/10 of the amount of maths you have ) and I have more common snese than you